OpenStudy (anonymous):

Find the Laplace transform Piecewise function: 2, 0<=x<3 -1, 3<=x

OpenStudy (anonymous):

Change it into the Heaviside unit step function.

OpenStudy (anonymous):

I guess i thought all you had to do was just treat it like a regular laplace transform problem but instead like for this one, integral from 0 to 3 of e^-sx * 2 dx + integral 3 to infinity of e^-sx * -1 dx.

OpenStudy (anonymous):

\[ \begin{split} F(x)&=\begin{cases} f(x)&x<c\\ g(x)&x>3 \end{cases}\\ &= f(x) + \begin{cases} 0&x<c\\ g(x)-f(x)&x>c \end{cases}\\ &= f(x) + (g(x)-f(x))\begin{cases} 0&x<c\\ 1&x>c \end{cases}\\ &= f(x) + (g(x)-f(x))\begin{cases} 0&x-c<0\\ 1&x-c>0 \end{cases}\\ F(x)&=f(x)+[g(x)-f(x)]H(x-c) \end{split} \]

OpenStudy (anonymous):

Yeah, I dunno my first instinct is to use Heaviside for some reason.

OpenStudy (anonymous):

well it talks about that in this section, but the examples he uses are really hard to follow... there isnt any specific examples on how to apply it to a problem like the one above, so there isn't much to follow.

OpenStudy (anonymous):

Well in this case \(f(x)=2\) and \(g(x)=-1\) So it would be \(2+(-1-2)H(x-c) = 2-3H(x-c)\)

OpenStudy (anonymous):

\[ \mathcal{L}\{2-3H(x-c)\}=\mathcal{L}\{2\}-3\mathcal{L}\{H(x-c)\} \]

OpenStudy (anonymous):

what is H(x-c)??

OpenStudy (anonymous):

\(H(x)\) is the standard Heaviside where the step is at \(x=0\). If you put in \(x-c\) then it shifts the step to be a \(x=c\)

OpenStudy (anonymous):

In general you want to remember that \[\large \mathcal{L}\{H(x-c)\} = \frac{e^{-cs}}{s} \]

OpenStudy (anonymous):

Okay I should have written: \[ 2+(-1-2)H(x-3) = 2-3H(x-3) \]Since the step is at \(x=3\)

OpenStudy (anonymous):

now when finding L[2] for the first part, are you just taking the integral of e^-sx * 2 and evaluating from 0 to infinity or are you evaluating from 0 to 3

OpenStudy (anonymous):

\[ \mathcal{L}\{2\} = 2\mathcal{L}\{1\} \]It's worth committing to memory that: \[ \mathcal{L}\{1\}=\frac{1}{s} \]

OpenStudy (anonymous):

okay so L[2] = 2/s, so then H[x-3] = e^-3x / s ??

OpenStudy (anonymous):

Yeah. but not \(x\) it's \(s\)

OpenStudy (anonymous):

sorry, yes. so then I will have (2/s) - (3e^-3s)/s = (2-3e^-3s)/s ?

OpenStudy (anonymous):

Yeah, I'm getting \[\Large \frac{2-3e^{-3s}}{s} \]

OpenStudy (anonymous):

That is what I got the first time as well., I dont have the answer but tried the same thing on a problem where the answer was given. 0, 0<=x<5 2, 5<=x<infinity what I got for my answer was 5 - [ (2+e^-5s / s] but the answer they got was 2e^-5s / s so I am not sure why the same doesnt work?

OpenStudy (anonymous):

Okay first of all change it from piece-wise to Heaviside.

OpenStudy (anonymous):

Use \[ \begin{split} F(x)&=\begin{cases} f(x)&x<c\\ g(x)&x>3 \end{cases}\\ &= f(x) + \begin{cases} 0&x<c\\ g(x)-f(x)&x>c \end{cases}\\ &= f(x) + (g(x)-f(x))\begin{cases} 0&x<c\\ 1&x>c \end{cases}\\ &= f(x) + (g(x)-f(x))\begin{cases} 0&x-c<0\\ 1&x-c>0 \end{cases}\\ F(x)&=f(x)+[g(x)-f(x)]H(x-c) \end{split} \] In this case \(f(x)=0,\;g(x)=5,\;c=5\)

OpenStudy (anonymous):

okay, i get it.. it just happened to work out the first time, but using that I get what the book gets

OpenStudy (anonymous):

but just for future problem, because I see them coming, if they have say x^2 for f(x) and say (x-5) for g(x), do you still use that same heaviside

OpenStudy (anonymous):

Yes, you still want to use the Heaviside.

OpenStudy (anonymous):

\[\Large \mathcal{L}\{H(x-c)f(x)\} = e^{-cs}\mathcal{L}\{f(x)\} \]It can only make things easier.

OpenStudy (anonymous):

thank you for your help... I appreciate all your time.